OpenStudy (anonymous):

Simplify. square root of 5 open parentheses 10 minus 4 square root of 2 close parentheses

OpenStudy (anonymous):

could you draw this?

OpenStudy (anonymous):

\[\sqrt{5}(10-4\sqrt{2})\] Is that what the problem is?

OpenStudy (anonymous):

14 15 square root of 2 5 square root of 2 end root minus 4 square root of 10 None of the above the options @wrstlr3232

OpenStudy (anonymous):

i believ its none of the above

OpenStudy (anonymous):

should be none of the above. is C \[5\sqrt{2}-4\sqrt{10}\]?